OpenStudy (anonymous):

what percent of 233,457 is 900; translate and solve

OpenStudy (anonymous):

divide 900 by 233,457 and convert the decimal to percent.

OpenStudy (nubeer):

(x/100) * 233457 = 900 find x from here.

OpenStudy (anonymous):

percent is out of one hundred, so if you want to find what percentage of 100 a certain number represents set up a ratio equation: \[\huge \frac{x}{100 \% }=\frac{900}{233457}\] read as: a certain number is to 100 percent as 900 is to 233457 Now moving to solve x you obtain: \[\huge x=\frac{900*100}{233457}[\%]\] Evaluate the number but also see how the units (percent) stick around